4.2 Part A Estimate the radiation pressure due to a bulb that emits 25 W of...

4.2

Part A

Estimate the radiation pressure due to a bulb that emits 25 W of EM radiation at a distance of 5.0 cm from the center of the bulb. Assume that light is completely absorbed. Express your answer to two significant figures and include the appropriate units.

Part B

Estimate the force exerted on a fingertip at this point. Suppose that the area of the fingertip is 1.4 cm2 . Assume that light is completely absorbed. Express your answer to two significant figures and include the appropriate units.
Answer #1

a)

Intensity is given by

I = P / (4 pi r^2)

I = 25 / ( 4 pi* 0.05^2)

I = 796.2 W/m^2

pressure is given by

Pressure = I/c = 2.654* 10^-6 N/m^2

=====

b)

F = Pressure * 1.4* 10^-4 = 3.715* 10^-10 N

=====

comment before rate in case any doubt, will reply for sure.. goodluck